OpenStudy (anonymous):

explain whether f(h(x)) and h(f(x)) will always result in the same number pleaseeeeeeeeee? I really need help and ill give out medals and become fans

OpenStudy (anonymous):

They ask you to create a new function, h(x). Then assign any number to x

OpenStudy (ipwnbunnies):

Just do an example. f(x) = x + 1 Let h(x) = x^2 Create both composite functions, see what you get.

OpenStudy (ipwnbunnies):

Hmm, that doesn't really make sense. o.o Create any function?

OpenStudy (anonymous):

here ill get the full question

OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (ipwnbunnies):

Ahh, ok. So, what was the function, f(x) that you created in problem 2?

OpenStudy (anonymous):

f(x) = 5x+2

OpenStudy (ipwnbunnies):

Ok, cool. Now, let's create use some simple function for h(x). Like h(x) = x^2

OpenStudy (anonymous):

ok

OpenStudy (ipwnbunnies):

Do you know what composite functions are? Do you understand what f(h(x)) and h(f(x)) mean?

OpenStudy (anonymous):

i understood them when they were asked in a question for, but not when i have to create them myself

OpenStudy (ipwnbunnies):

Oh, it's easier when you have to create them yourself. XD It makes the question simpler. So, in that case, can you find f(h(x))?

OpenStudy (anonymous):

so, f is 5 + 2

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

5x*

OpenStudy (ipwnbunnies):

f(x) = 5x + 2 And for f(h(x)), we want to plug in h(x) into each x of f(x). h(x) = x^2 f(h(x)) = 5(x^2) + 2

OpenStudy (ipwnbunnies):

That simple. Do the similar thing to h(f(x)). And you'll see if they're similar or not.

OpenStudy (anonymous):

but what is x?

OpenStudy (anonymous):

i dont understand how to plug in

OpenStudy (ipwnbunnies):

I think they were saying just to use 'x' as a variable for a number.

OpenStudy (ipwnbunnies):

h(x) = x^2 For h(f(x)), you're plugging in f(x) into the 'x' in h(x).

OpenStudy (anonymous):

so i plug in 5x +2 ?

OpenStudy (ipwnbunnies):

Yes, into the x in x^2.

OpenStudy (anonymous):

how would that work im so confused sorry

OpenStudy (anonymous):

oh ok so 5^2 + 2?

OpenStudy (ipwnbunnies):

h(f(x)) = (5x + 2)^2 I just plugged the entire function of f(x) into x. And the entire thing is squared.

OpenStudy (anonymous):

but how do you solve that?

OpenStudy (ipwnbunnies):

You can expand it out. We're just making functions, not evaluating them. They ask you to determine whether f(h(x)) and h(f(x)) will give you the same result.

OpenStudy (anonymous):

so you don't need to solve it?

OpenStudy (anonymous):

so for f(h(x)) it would be 5x+2(^2) ???

OpenStudy (ipwnbunnies):

Nooo.f(h(x)), we plugged h(x) into the x in f(x).